OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)?

OpenStudy (marziajackson$):

what are the choice's?

OpenStudy (anonymous):

4x – 2y = –6 y = 2x – 5. y = 3x + 4 2x + y = –1

OpenStudy (anonymous):

@MarziaJackson$

OpenStudy (marziajackson$):

I going to say the first choice